a man is 37 years old and his child's age is 8,how many years ago was the
product of their age at 69

Answer by ankor@dixie-net.com(22740)

a man is 37 years old and his child's age is 8,how many years ago was the
product of their age at 69
the way this is written will not give an integer solution. Perhaps the 69 should be 96
:
let y = no. of years for this to be true
(37-y)*(8-y) = 96
FOIL
296 - 37y - 8y + y^2 = 96
Arrange to form a quadratic equation
y^2 - 45y + 296 - 96 = 0
y^2 - 45y + 200 = 0
Factors to
(x-40)(x-5) = 0
two solutions
x = 40, obviously not possible
and
x = 5 yrs ago the product of their ages will be 96
:
:
Check; 5 yrs ago the man was 32 and the child was 3
3 * 32 = 96